# A Visual Perception-Based Tunable Framework and Evaluation Benchmark for H.265/HEVC ROI Encryption

### v1

@@ Abstract
ROI selective encryption, as an efficient privacy protection technique, encrypts only the key regions in the video, thereby ensuring security while minimizing the impact on coding efficiency. However, existing ROI-based video encryption methods suffer from insufficient flexibility and lack of a unified evaluation system. To address these issues, we propose a visual perception-based tunable framework and evaluation benchmark for H.265/HEVC ROI encryption. Our scheme introduces three key contributions: 1) A ROI region recognition module based on visual perception network is proposed to accurately identify the ROI region in videos. 2) A three-level tunable encryption strategy is implemented while balancing security and real-time performance. 3) A unified ROI encryption evaluation benchmark is developed to provide a standardized quantitative platform for subsequent research. This triple strategy provides new solution and significant unified performance evaluation methods for ROI selective encryption field. Experimental results indicate that the proposed benchmark can comprehensively measure the performance of the ROI selective encryption. Compared to existing ROI encryption algorithms, our proposed enhanced and advanced level encryption exhibit superior performance in multiple performance metrics. In general, the proposed framework effectively meets the privacy protection requirements in H.265/HEVC and provides a reliable solution for secure and efficient processing of sensitive video content.
@@ I Introduction
Therefore, protecting sensitive information while maintaining the usability of video data for automated analysis tasks (such as behavior recognition) remains a significant challenge. To address this issue, researchers have proposed Region of Interest (ROI) encryption. The core concept of ROI encryption is to selectively encrypt only sensitive regions within video frames while keeping non-sensitive areas unencrypted. Early naive ROI encryption methods performed encryption directly on raw pixel data prior to compression, severely disrupting the spatial correlations within the video [carrillo2008compression, dufaux2008h, dufaux2008scrambling, hosny2022privacy, li2024ppl]. As a result, the compression efficiency of subsequent codecs (e.g., H.265/HEVC) was significantly reduced, and bitstream compatibility could no longer be guaranteed [stutz2011survey].
@@
Consequently, Region of Interest (ROI) selective encryption (SE) [xu2013roi], a strategy that integrates encryption with video encoding, has become a key research focus in the field of video security [sheng2024content]. Since H.265/HEVC is currently one of the most widely adopted video coding standards [van2013encryption], ROI SE techniques based on H.265/HEVC have become a major research focus in the field of video encryption [peng2013roi, im2024cabac, yu2023coding, taha2018end, sohn2009privacy, farajallah2015roi]. Such approaches preserve bitstream format compliance with minimal bitrate overhead and enable standard decoders to reconstruct non-ROI regions without requiring access to the encryption key. Despite the great potential of ROI SE based on H.265/HEVC, existing algorithms still face two major challenges: (1) Lack of flexibility in encryption schemes: existing methods use a fixed combination of syntax elements for encrypting ROI regions. However, as shown in Fig. 1, different application scenarios exhibit diverse requirements for security and real-time performance. Therefore, existing methods struggles to meet diverse needs and fails to achieve a dynamic balance between security and real-time performance. (2) Lack of a Unified and Standardized Evaluation Benchmark: This remains the most critical issue hindering progress in this field. We reviewed some typical ROI encryption studies based on H.264/AVC and H.265/HEVC, and summarized their evaluation systems in Table I. As shown in the table, there is considerable variation in the performance metrics adopted across different studies. Some works primarily focus on efficiency-related indicators, such as encoding time [peng2013roi, yu2023coding], while others emphasize visual quality and bit rate change[taha2018end, farajallah2015roi].
TABLE I: Summary of evaluation metrics in some typical ROI SE schemes
Author Standard Experimental Focus
@@
Peng et al. [peng2013roi] H.264/AVC Key security, Encoding time, Bit rate change
@@
Yu et al. [yu2023coding] H.265/HEVC PSNR, SSIM, IoU, Encoding time
@@
Taha et al. [taha2018end] H.265/HEVC PSNR, SSIM, Bit rate change, Key security
Sohn et al. [sohn2009privacy] H.264/SVC Bit rate change
@@
Farajallah et al. [farajallah2015roi] H.265/HEVC Bit rate change
@@
This arbitrariness in metric selection, combined with the lack of transparency in testing methodologies, has led to the absence of a unified, comprehensive, and quantifiable benchmarking framework in the field. As a result, the primary issue to be addressed is building a standard evaluation benchmark. Based on this, we proposea visual perception-based tunable framework and evaluation benchmark for H.265/HEVC ROI encryption. Firstly, a visual perception network is used to identify ROI coordinates and segment the video frame into ROI and non-ROI region. Then, a three-level tunable ROI encryption strategy (the basic level, the enhanced level, and the advanced level) is applied to encrypt ROI region of the video based on the user requirements. Finally, a unified ROI encryption evaluation benchmark is developed to better measure the performance of ROI encryption. In summary, the main contributions of this paper are as follows:
@@ II-A ROI Naive Encryption
Although ROI naive encryption can protect privacy by directly encrypting sensitive regions, it faces several technical challenges of a significant decrease in video coding efficiency, with bit rate increases ranging from 8% to 23% , Therefore, ROI selective encryption is proposed to solve the problems.
@@ III-A Object Detection Based on Visual Perception Network
The syntax elements is the basic units used to represent video information during the encoding process. In H.265/HEVC, the video frame is compressed into a bitstream after prediction, transformation, quantization and entropy coding. In this process, a large number of syntax elements are obtained. TABLE II provides the main syntax elements in H.265/HEVC. These syntax elements are binarized using three methods: $k$-order Exponential Golomb ($\mathrm{EG}_{k}$) coding, $k$-order Truncated Rice ($\mathrm{TR}_{k}$) coding, and Fixed-Length (FL) coding[flynn2015overview]. After the binarization stage, CABAC is applied. H.265/HEVC has two entropy coding modes in CABAC: regular mode and bypass mode. In the regular mode, the encoding of each binary value is based on an adaptively updated probability model. In the bypass mode, the encoding of syntax elements is based on fixed probabilities[marpe2010entropy].
@@
While encrypting the syntax elements in regular mode will inevitably result in a change in the bit rate, and encrypting syntax elements in bypass mode will not increase the bit rate [liu2010survey]. Based on this, we analyze the impact of encryption of each syntax element below.
As shown in TABLE II, syntax elements are divided into three categories: prediction, residual, and filtering. In video coding, decoded pixels are primarily restored by prediction and residual data. Thus, encrypting prediction and residual syntax elements will cause a strong perturbation effect[stutz2011survey]. Furthermore, according to the analysis in [tang2022format], SAO parameters do not significantly contribute to visual distortion. To enhance efficiency, it is recommended not to encrypt these parameters.
In conclusion, encrypting IPM, MVD, Residual, and Delta QP offers a higher scrambling effect, and encrypting IPM, Merge index, MVPIdx, RefFrmIdx, and Delta QP inevitably change the bit rate. Additionally, the encryption must strictly adhere to the encoding rules of syntax elements to ensure that the encrypted bitstream remains compatible with the H.265/HEVC format.
@@
Based on the characteristics of syntax elements, we propose a graded encryption strategy for H.265/HEVC with three levels of encryption, including basic level, enhanced level, and advanced level. This strategy allows for flexible selection of different encryption methods for ROI regions according to specific needs. The complete framework of the proposed three-level encryption strategy is shown in Fig. 3, which consists of two core components: ROI region recognition module and three-level tunable ROI selective encryption.
Figure 3: The proposed three-level encryption strategy
@@ IV-B1 Basic Level Encryption
The basic level encryption aims to ensure real-time performance and zero bit increment during encoding, while providing basic visual distortion effects. It is suitable for scenarios with high real-time requirements. This strategy involves encrypting certain syntax elements such as MVD sign and value, residual sign and value, the suffix of Delta QP, Merge index in bypass mode, and the suffix of RefFrmIdx in bypass mode. The specific encryption methods for these syntax elements are described below.
@@ IV-B3 Advanced Level Encryption
The edge region contains the highest concentration of video frame information, making its protection essential [zhang2013edge]. Existing ROI encryption schemes typically perturb syntax elements but neglect edge information. To address this, we proposed an improved coefficient scrambling algorithm based on [peng2019tunable] for edge region in the advanced level encryption, building upon the enhanced level encryption. The brief steps of the coefficient scrambling algorithm are as follows:
Step 1: The Canny operator $Canny(\cdot)$ is applied to filter the luma component of a frame $FL$ to obtain the pixel edge information $FL^{edge}(x,y)$:
$FL^{edge}(x,y)=Canny(FL(x,y)).$ (22)
@@
Step 5: Reversibly embed the sign into the last non-zero scanned coefficient $LastCoef$ in the current TU, as shown below:
@@ V ROI Encryption Evaluation Benchmark
ROI encryption have recently garnered significant attention in the field of video encryption. However, the current research on ROI encryption often lacks comprehensive experimental evaluation, and the measurement indicators used are not standardized. For instance, Peng[peng2013roi] assessed encoding time and compression ratio. Yu[yu2023coding] focused on testing the IoU between the encrypted region and the ground truth, as well as encryption time. Taha[taha2018end] evaluated their scheme by testing bit rate change, PSNR and SSIM values. These evaluation methods are relatively simplistic, lacking sufficient persuasive power to demonstrate the superiority of the scheme. Furthermore, the inconsistency in measurement indicators complicates effective horizontal comparisons and comprehensive evaluations between different schemes.

### v2

@@ Abstract
ROI selective encryption, as an efficient privacy protection technique, encrypts only the key regions in the video, thereby ensuring security while minimizing the impact on coding efficiency. However, existing ROI-based video encryption methods suffer from insufficient flexibility and lack of a unified evaluation system. In particular, the lack of standard testing platforms is a major obstacle to the development of this field. To address these issues, we propose a visual perception-based tunable framework and evaluation benchmark for H.265/HEVC ROI encryption. Our scheme introduces three key contributions: 1) A ROI region recognition module based on visual perception network is proposed to accurately identify the ROI region in videos. 2) A three-level tunable encryption strategy is implemented while balancing security and real-time performance. 3) A unified ROI encryption evaluation benchmark is developed to provide a standardized quantitative platform for subsequent research. This triple strategy provides new solution and significant unified performance evaluation methods for ROI selective encryption field. Experimental results indicate that the proposed benchmark can comprehensively measure the performance of ROI selective encryption and provide a complete testing solution for this field. Compared to existing ROI encryption algorithms, our proposed enhanced and advanced level encryption exhibit superior performance in multiple performance metrics. In general, the proposed framework effectively meets the privacy protection requirements in H.265/HEVC and provides a reliable solution for secure and efficient processing of sensitive video content.
@@ I Introduction
Therefore, protecting sensitive information while maintaining the usability of video data for automated analysis tasks (such as behavior recognition) remains a significant challenge. To address this issue, researchers have proposed Region of Interest (ROI) encryption. The core concept of ROI encryption is to selectively encrypt only sensitive regions within video frames while keeping non-sensitive areas unencrypted. Early naive ROI encryption methods performed encryption directly on raw pixel data prior to compression, severely disrupting spatial correlations within the video [carrillo2008compression, dufaux2008h, dufaux2008scrambling, hosny2022privacy, li2024ppl]. As a result, the compression efficiency of subsequent codecs (e.g., H.265/HEVC) was significantly reduced, and bitstream compatibility could no longer be guaranteed [stutz2011survey].
@@
Consequently, Region of Interest (ROI) Selective Encryption (SE) [xu2013roi], a strategy that integrates encryption with video encoding, has become a key research focus in the field of video security [sheng2024content]. Since H.265/HEVC is currently one of the most widely adopted video coding standards [van2013encryption], ROI SE techniques based on H.265/HEVC have become a major research focus in the field of video encryption [peng2013roi, im2024cabac, yu2023coding, taha2018end, sohn2009privacy, farajallah2015roi]. Such approaches preserve bitstream format compliance with minimal bitrate overhead and enable standard decoders to reconstruct non-ROI regions without requiring access to the encryption key. Despite the great potential of ROI SE based on H.265/HEVC, existing algorithms still face two major challenges: (1) Lack of Flexibility in Encryption Schemes: Existing methods typically employ a fixed combination of syntax elements when encrypting ROI regions. For example, Peng et al. [peng2013roi] encrypts CAVLC, IPM, and MVD, achieving relatively high security but introducing approximately an 8% increase in bitrate. In contrast, Sohn et al. [sohn2009privacy] only encrypts the transform quantization coefficients, resulting in a low bit rate increase. However, the limited distortion leads to insufficient overall security. These single encryption strategies are difficult to apply in complex real-world scenarios, since different application scenarios exhibit diverse requirements for security and real-time performance, which can be shown in Fig. 1. Therefore, existing methods fail to achieve a dynamic balance between security and real-time performance. (2) Lack of a Unified and Standardized Evaluation Benchmark: The absence of a unified evaluation benchmark remains a key challenge in the ROI encryption field. TABLE I summarizes several representative ROI encryption studies based on H.264/AVC and H.265/HEVC, including their datasets, experimental focus, and comparison algorithm. As shown in TABLE I, the selected datasets of these studies vary widely, ranging from 10 QCIF sequences with unspecified sources to standardized test sets such as Xiph.org, JCT-VC common test conditions, and SPEVI. Moreover, there is substantial diversity in the performance metrics, some studies primarily consider efficiency-related indicators, such as encoding time [peng2013roi, yu2023coding], while others focus on visual quality, bit rate change, and key security [taha2018end, sohn2009privacy, farajallah2015roi]. In addition, we can also find from the table that the comparison algorithms are also inconsistent, with most studies lacking direct comparisons, except for a few that reference prior methods.
TABLE I: Summary of some typical ROI SE schemes
Scheme Standard Dataset Experimental focus Comparison algorithm
@@
Peng et al. [peng2013roi] H.264/AVC 10 QCIF sequences (unspecified source) Key security, Encoding time, Bit rate change None
@@
Yu et al. [yu2023coding] H.265/HEVC Xiph.org video test media PSNR, SSIM, IoU, Encoding time None
@@
Taha et al. [taha2018end] H.265/HEVC JCT-VC common test conditions PSNR, SSIM, Bit rate change, Key security Van et al. [van2013encryption], Boyadjis et al. [boyadjis2016extended]
Sohn et al. [sohn2009privacy] H.264/SVC SPEVI dataset Bit rate change None
@@
Farajallah et al. [farajallah2015roi] H.265/HEVC JCT-VC common test conditions Bit rate change None
@@
This arbitrariness in metric selection, combined with the lack of transparency in testing methodologies, has led to the absence of a unified, comprehensive, and quantifiable benchmark in the field. Therefore, we propose a visual perception-based tunable framework and evaluation benchmark for H.265/HEVC ROI encryption. Firstly, a visual perception network is used to identify ROI coordinates and segment the video frame into ROI and non-ROI region. Then, a three-level tunable ROI encryption strategy (the basic, enhanced, and advanced level) is applied to encrypt ROI region of the video based on the user requirements. Finally, a unified ROI encryption evaluation benchmark is developed to better measure the performance of ROI encryption. We hope that the proposed benchmark can provide a reference testing platform for this field. In summary, the main contributions of this paper are as follows:
@@ II-A ROI Naive Encryption
Although ROI naive encryption can protect privacy by directly encrypting sensitive regions, it faces the technical challenges of a significant decrease in video coding efficiency, with bit rate increases ranging from 8% to 23% , Therefore, ROI selective encryption is proposed to solve the problems.
@@ III-A Object Detection Based on Visual Perception Network
The syntax elements is the basic units used to represent video information during the encoding process. In H.265/HEVC, the video frame is compressed into a bitstream after prediction, transformation, quantization and entropy coding. In this process, a large number of syntax elements are obtained. TABLE II provides the main syntax elements in H.265/HEVC. These syntax elements are binarized using three methods: $k$-order Exponential Golomb ($\mathrm{EG}_{k}$) coding, $k$-order Truncated Rice ($\mathrm{TR}_{k}$) coding, and Fixed-Length (FL) coding[flynn2015overview]. After the binarization stage, CABAC is applied. H.265/HEVC has two entropy coding modes in CABAC: regular mode and bypass mode. In the regular mode, the encoding of each binary value is based on an adaptively updated probability model. In the bypass mode, the encoding of syntax elements is based on fixed probabilities.
@@
While encrypting the syntax elements in regular mode will inevitably result in a change in the bit rate, and encrypting syntax elements in bypass mode will not increase the bit rate [liu2010survey]. Based on this, we analyze the impact of encryption of each syntax element. As shown in TABLE II, syntax elements are divided into three categories: prediction, residual, and filtering. In video coding, decoded pixels are primarily restored by prediction and residual data. Thus, encrypting prediction and residual syntax elements will cause a strong perturbation effect[stutz2011survey]. Furthermore, according to the analysis in [tang2022format], SAO parameters do not significantly contribute to visual distortion. To enhance efficiency, it is recommended not to encrypt these parameters. In conclusion, encrypting IPM, MVD, Residual, and Delta QP offers a higher scrambling effect, and encrypting IPM, Merge index, MVPIdx, RefFrmIdx, and Delta QP inevitably change the bit rate. Additionally, the encryption must strictly adhere to the encoding rules of syntax elements to ensure that the encrypted bitstream remains compatible with the H.265/HEVC format.
@@
Based on the characteristics of syntax elements, we propose a graded encryption strategy for H.265/HEVC with three levels of encryption, including basic level, enhanced level, and advanced level. This strategy allows for the flexible selection of different encryption methods based on specific needs. The complete framework of the proposed three-level encryption strategy is shown in Fig. 3. In the ROI region recognition module, a visual perception network first extracts the ROI coordinates from each input video frame. These coordinates are then converted into tile-level ROI masks. Finally, based on specific application requirements, the proposed three-level tunable ROI selective encryption strategy is subsequently applied to encrypt the syntax elements within the marked tile-level ROI regions. Therefore, our encryption framework consists of two core components: ROI region recognition module and three-level tunable ROI selective encryption.
Figure 3: The framework of the proposed encryption strategy
@@ IV-B1 Basic Level Encryption
The basic level encryption aims to ensure real-time performance and zero bit increment, while providing basic visual distortion effects. It is suitable for scenarios with high real-time requirements. This strategy involves encrypting certain syntax elements such as MVD sign and value, residual sign and value, the suffix of Delta QP, Merge index, and the suffix of RefFrmIdx in bypass mode.
@@ IV-B3 Advanced Level Encryption
The edge contains the highest concentration of video information, making its protection essential [zhang2013edge]. Existing ROI encryption typically perturbs syntax elements but neglects edge information. To address this, we proposed an improved coefficient scrambling algorithm based on [peng2019tunable] for the edge region in the advanced level encryption, building on the enhanced level encryption. The brief steps of the coefficient scrambling algorithm are as follows:
Step 1: The Canny operator $Canny(\cdot)$ is applied to filter the luma component of a frame $FL$ to obtain the pixel edge information $FL^{\text{edge}}(x,y)$:
$FL^{\text{edge}}(x,y)=Canny(FL(x,y)).$ (22)
@@
Step 5: Reversibly embed the sign into the last non-zero scanned coefficient $LastCoef$ in the current TU:
@@ V ROI Encryption Evaluation Benchmark
ROI encryption have recently garnered significant attention in the field of video encryption. However, the current research on ROI encryption often lacks comprehensive experimental evaluation, and the dataset, comparison algorithm, and evaluation methods used are not standardized. For instance, Peng[peng2013roi] assessed encoding time and compression ratio with no comparison algorithm in 10 QCIF sequences with unspecified sources. Yu[yu2023coding] focused on testing IoU between the encrypted region and the ground truth, as well as encryption time, but did not include any comparisons on the Xiph.org dataset. Taha[taha2018end] evaluated their scheme by testing bit rate change, PSNR and SSIM values with Van et al. [van2013encryption] and Boyadjis et al. [boyadjis2016extended] in the JCT-VC common test dataset. These evaluation methods are relatively simplistic, lacking sufficient persuasive power to demonstrate the superiority of the scheme. Furthermore, the inconsistency in dataset, comparison algorithm, and evaluation methods complicates effective horizontal comparisons and comprehensive evaluations between different schemes.
In conclusion, this field is in urgent need of a standardized evaluation benchmark. Therefore, we pioneer a new comprehensive ROI encryption evaluation benchmark to provide a standardized evaluation platform for different ROI encryption algorithms. We believe that the performance comparison of an ROI encryption algorithm depends on three key factors: datasets, comparison algorithm, and evaluation methods. (All the contents of our proposed benchmark are available at https://github.com/playboiwg/ROI-encryption-evaluation-benchmark/tree/main)
V-A Datasets
Existing ROI encryption studies exhibit considerable arbitrariness in dataset selection, often employing QCIF sequences of unclear origin or outdated formats, alongside diverse test sets from sources such as Xiph.org, JCT-VC, and SPEVI. This lack of unified standards severely impedes fair, comprehensive, and quantifiable cross-comparisons within the field. To address this issue, a comprehensive evaluation benchmark should first establish a standardized set of video datasets, ensuring that different encryption schemes can be objectively assessed and compared under identical testing conditions. Although there are many publicly available video datasets, not all of them contain ROI regions. Therefore, we compile and collect totally 55 YUV sequences that contain ROI regions from widely recognized public datasets commonly used in video coding and video processing, which can serve as standard test sets for future studies. TABLE III summarizes these public datasets and their main characteristics. In the future, we will also further collect and expand our ROI encryption dataset.
TABLE III: Standard video datasets constructed for evaluation
Dataset Resolution Quantity Link
UVG Dataset 4K 7 https://ultravideo.fi/dataset.html
Xiph.org Media QCIF – 4K 18 https://media.xiph.org/video/derf/
MSU CVQAD 240P – 4K 30 https://videoprocessing.ai/datasets/cvqad.html
UVG Dataset. We selected 7 sequences containing ROIs from the UVG dataset. These sequences were originally captured at 120 fps with a resolution of $3840\times 2160$. They contain clear human subjects and specific motion patterns, such as Beauty (close-up face) and Jockey (high-speed human motion). These videos possess extremely high texture details, enabling the simulation of privacy protection requirements in ultra-clear surveillance or film production scenarios.
Xiph.org Media. We constructed a test set of 18 sequences containing ROIs from the Xiph.org repository, with resolutions ranging from $176\times 144$ to $3840\times 2160$ and frame rates ranging from 25 to 60 fps. Focusing on typical application scenarios such as video conferencing, news broadcasting, and traffic surveillance, this selection covers diverse ROIs ranging from single faces (e.g., Foreman, Akiyo) to multi-target pedestrians and vehicles (e.g., Bus, Coastguard).
MSU CVQAD. We incorporated 30 video clips containing ROIs from the MSU CVQAD database, with resolutions ranging from $480\times 240$ to $3840\times 2160$ and a frame rate of 30 fps. We specifically selected videos featuring high-density crowds and complex street backgrounds. These scenes are often accompanied by illumination changes and object occlusions. Testing on these sequences allows evaluation of the algorithm’s encryption performance in complex environments, simulating real urban security surveillance scenarios.
V-B Comparison Algorithms
To enable a comprehensive and unified evaluation, our benchmark incorporates six representative ROI selective encryption schemes reported in the existing literature, covering multiple coding standards such as H.264/AVC, H.265/HEVC, and H.266/VVC. All representative algorithms are summarized and analyzed to establish a more complete comparison framework. The implementation details and typical advantages and limitations of each method are presented in TABLE IV, and their brief introductions are as follows:
TABLE IV: Summary of representative ROI SE schemes
Scheme Standard Implementation detail Advantage Limitation
Peng et al. [peng2013roi]. H.264/AVC Uses FMO to separate ROI and non-ROI, encrypts IPM and MVD symbol bits using a chaotic system High format compatibility and strong security High bitrate increase
Kim et al. [kim2007selective] H.264/SVC Uses PRNG-generated sequence to encrypt MVD within ROI High real-time performance Limited security
Yu et al. [yu2023coding] H.265/HEVC Encrypts IPM, MV symbol bits, and TC symbol bits using AES-CFB generated keys Fine-grained encryption, high IoU accuracy High bitrate increase
Taha et al. [taha2018end] H.265/HEVC Encrypts IPM, MVD, MV symbol bits, and TC using a chaotic system Good real-time performance High bitrate increase
Im et al. [im2024cabac] H.266/VVC Encrypts MVD and residual signals using AES-128 Good real-time performance Possible error propagation
Im et al. [im2023roi] H.266/VVC Encrypts MVD, residual signals, and QP using AES-128 Higher security Reduced real-time performance
Peng et al. [peng2013roi] This comparative algorithm is designed for H.264/AVC, it employs FMO to isolate ROI macroblocks into independent slice groups and employs a chaotic system to encrypt the IPM and MVD.
Kim et al. [kim2007selective] This comparative algorithm is designed for H.264/SVC, it uses a sequence generated by a pseudo-eandom number generator to selectively encrypt the MVD within the ROI layers.
Yu et al. [yu2023coding] This comparative algorithm is designed for H.265/HEVC, it utilizes the AES-CFB mode to generate keystreams for the joint encryption of IPM, MV sign bits, and TC sign bits.
Taha et al. [taha2018end] This comparative algorithm is based on the H.265/HEVC standard, which employs a chaotic system to encrypt IPM, MVD, MV sign bits, and TC within the ROI.
Im et al. [im2024cabac] This comparative algorithm is designed for the H.266/VVC standard, and it uses the AES-128 algorithm within the CABAC entropy coding process to encrypt MVD and residual signals in the ROI.
Im et al. [im2023roi] This algorithm serves as an extension of [im2024cabac], further utilizing the AES-128 algorithm to encrypt MVD, residual signals, and QP to enhance encryption strength.
To ensure fairness and standardization in comparisons, we recommend selecting the corresponding algorithms from the aforementioned representative schemes as comparison baselines according to the adopted video coding standard.
V-C Evaluation Methods